A and B are the point (3,-1) and (2,3) respectively.
L is the line 8y - 2x - 3 = 0
show that the line L is perpendicular to AB

Answer by richwmiller(17219)

y = 11-4x
-4 is the slope
8y - 2x - 3 = 0
y = x/4+3/8
1/4 is the slope
1/4*-4=-1
they are perpendicular.